OpenStudy (anonymous):

w=2wxyz solve for y Please Help!!

OpenStudy (anonymous):

divide both sides by w: 1 = 2xyz divide both sides by 2xz: 1/(2xz) = y

OpenStudy (anonymous):

Divide everything to isolate y. w/2wxz = y. w on the top and bottom can be cancelled. y=1/2xz

OpenStudy (anonymous):

Thank you, Thank you!!